OpenStudy (goformit100):

Find the last two digits of (right most) of 3^999.

OpenStudy (goformit100):

@cinar @noided

OpenStudy (anonymous):

67

OpenStudy (anonymous):

3^999=440690273160268878963485086584048121988474010917382722554973456075609532448901633180259437950202687321303259232290860785316984860700206303955114241752651224675873408399440267959338258076321613758130133372529539347042982605207698146020522057684695558163502059375160114801849018132346298605821789418305378740276756187926194096742805466102629298972852134694966312536457747390615453312898505588339646862703020142029890479621367604783461882915721944003538122044057700922967618406667

OpenStudy (goformit100):

Thank you Sir

OpenStudy (nathan917):

I'm going to use modulo notation. We need to find 3^999 (mod 100) By Euler's Theorem, since gcd(3, 100) = 1 we see that 3^(phi(100)) = 1 (mod 100) phi(100) = phi(2^2 * 5^2) = phi(2^2) * phi(5^2) = 2^1 (2 - 1) * 5^1 (5 - 1) phi(100) = 2 * 20 = 40 Thus, 3^40 = 1 (mod 100). Let 3^999 = x (mod 100). Multiply both sides by 3. Then 3^1000 = 3x (mod 100) (3^40)^25 = 3x (mod 100) 1^25 = 3x (mod 100) 1 = 3x (mod 100) Note that if x = -33, then 1 = -99 (mod 100) which is true since 100 | (1 - -99). Note that -33 = 67 (mod 100) since 100 | (-33 - 67). Answer: 67

OpenStudy (goformit100):

Thank you Sirr